OpenStudy (anonymous):

find the area of a triangle where the base (b) is 12 centimeters and the height (h) is 9 centimeters?

OpenStudy (lgbasallote):

hint: formula for triangle is \[A = \frac{bh}2\] does that help?

OpenStudy (anonymous):

(9*12)/2 =108/2 =54 sq. cm

OpenStudy (anonymous):

\[\frac{ bh }{ 2 } = \frac{ 9 \times 12 }{ 2 }\]